Circular No. 7909

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2002 J6, C/2002 J7, C/2002 K3 (SOHO) Further to IAUC 7899, D. Hammer reports his measurements for ad-

ditional Kreutz sungrazing comets found by X. Leprette (C/2002 J6), R. Matson (C/2002 J7), and R. Kracht (C/2002 K3) on SOHO website C2 images. The reduced observations and orbital elements by B. G. Marsden are included on MPEC 2002-K48.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 J6 May 14.701  3^{h}29.^{m}6  +17^{o}19′
C/2002 J7  15.963  3 34.0  +17 33
C/2002 K3  18.671  3 44.3  +18 09

COMET 2002 K4 S. Pravdo, E. Helin, and K. Lawrence, Jet Propulsion Laboratory,

report the discovery of a comet (with nuclear condensation of diameter ∼ 4′′ and tail extending ∼ 10′′ to the southwest) on NEAT images taken at Haleakala on May 27.5 UT, with confirming NEAT images taken at Palomar on May 28.

2002 UT  α_{2000}  δ_{2000}  m_{1}
May 27.50576  21^{h}46^{m}31.^{s}27  −0^{o}16′08′′.5  18.5
27.51644  21 46 30.83  −0 15 53.8  18.8
27.52693  21 46 30.42  −0 15 40.3  18.5
28.44205  21 45 52.78  +0 04 35.3  17.2
28.45276  21 45 52.32  +0 04 49.9  17.3
28.46348  21 45 51.84  +0 05 04.4  17.3

COMET C/2002 F1 (UTSUNOMIYA)

Ephemeris extension to IAUC 7888 (elements MPC 45655):

2002TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 26 5^{h}43.^{m}34 + 4^{o}22.^{′}0 1.657 0.900 28.^{o}1 32.^{o}0  9.1
31 5 57.52 + 1 00.3 1.736 0.992 29.7 30.4  9.7
June 5 6 10.44 − 2 03.9 1.813 1.082 31.3 29.1 10.1
10 6 22.40 − 4 53.8 1.889 1.172 32.8 28.0 10.6
15 6 33.63 − 7 32.2 1.963 1.260 34.2 27.0 11.0
20 6 44.26 −10 01.1 2.035 1.347 35.7 26.1 11.3
25 6 54.41 −12 22.5 2.105 1.432 37.1 25.3 11.7
2002 May 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green